OpenStudy (anonymous):

16. Jesse played two days of golf. On the second day, he got a score of 2 below par, or -2. His total score for the two days was 8 below par or -8. What was Jesse's score the first day? a) Choose and label a variable for this situation. b) Write an equation using the variable that would answer the question. c) Solve your equation, and check your answer. d) Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (espex):

Well you first want to lay out what you know: Total: -8 Day 2: -2 Day 1: ? Now you know that both days together equals the total, so D1+D2=Tot, so the next thing to do is look at your questions and see what you can use to start filling in. a) Choose a variable, well the most common variable is 'x', which you will put in the problem for what you don't know--Day 1. b) Write an equation that will answer this problem, since the score on day 1 plus the score on day 2 equals the total score then you can write, x+(-2)=-8 c)Solve your equation,\[ x+(-2)=-8 \rightarrow x=-8+2 \rightarrow x=-6\] then test your solution, -6+-2=-8 (check!) d)Now put your solution into a "word solution" for the "word problem", Jesse played two days of golf, on the first day he got 6 under par, or -6, on the second day he got 2 under par, or -2, his total score for the two days was 8 under par, or -8.